yomomalikebracobama:

Solve y over negative 6 + 5 = 9. (1 point) a 24 b −24 c 78 d −78

axie:

i got this @Astro

yomomalikebracobama:

yay

axie:

wait what re word the question...

axie:

OHHH nvm i got it

yomomalikebracobama:

one sec it dontshow the fraction

axie:

so plug in each number for the question

axie:

so like 24/6+5=?

axie:

-24/6+5=?

yomomalikebracobama:

no

axie:

wdym no?

yomomalikebracobama:

y/6+5=9

yomomalikebracobama:

-6 *

axie:

yes i know

yomomalikebracobama:

axie:

you have to plug in each number to figure out which number equals 9

axie:

yes plug in the answer choices for y to figure out your fianl answer

yomomalikebracobama:

-24?

axie:

so it u plug in -24 or 24 this would be your final snwer

axie:

@axie wrote:
so it u plug in -24 or 24 this would be your final snwer

1 attachment
axie:

waittt

axie:

hold on

yomomalikebracobama:

no-24 is da awnser u said plug it in so i put the prblem into google

yomomalikebracobama:

and itgave me the awnser

axie:

actually yes you are right sorry.

axie:

24/6 equals 4 plus 5 equals 9

yomomalikebracobama:

GOOGLE IS RIGHT DONT GET ME ARSTED FOR TAKING CREDIT